OpenStudy (anonymous):

im in community college math 107 the question asks determine which values could be used to clear the frections in the given equation. select all that apply. 1/3(x) - 5/6 = x/9 a. 6 b. 9 c. 12 d. 18 e. 33 f. 36 right off the practice quiz

OpenStudy (anonymous):

18 and 36

OpenStudy (anonymous):

yea but howd you get that...

OpenStudy (anonymous):

figure out which numbers 3,6 and 9 go into without any remainders.

OpenStudy (anonymous):

ohhh i seee now thank you

OpenStudy (anonymous):

no prob :)

OpenStudy (anonymous):

okay so i distributed 18/1(1/3x-5/6) = x/9(18/1) and got 6x-15=2x -15= -4x 3.75=x.... i dont get how thats what x is...

OpenStudy (anonymous):

is the question just asking to find what numbers the denominators can go into?